Circular No. 6914

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cf IN NGC 3504 Apparently independent discovery reports of a supernova in NGC 3504

have been received (both from observations made at Mauna Kea) from Eric Emsellem, Observatoire de Lyon, via CCD images obtained on Mar. 31.272 UT at the Canada-France-Hawaii Telescope (CFHT), and from Geoffrey Clayton and Karl Gordon, Louisiana State University, via separate J, H, and K images taken on Mar. 2.38 with the NASA Infrared Telescope Facility. Emsellem notes that the star is located at α = 11^{h}03^{m}11^{s}.21, δ = +27^{o}58′21′′.0 (equinox 2000.0), which is 0′′.6 east and 5′′.3 south of the center of NGC 3504. Clayton provides position end figures 12^{s}.01, 13′′.2. Emsellem calculates that the object had a flux of 6.7 × 10^{−16} erg s^{−1} cm^{−2} Å^{−1} at 900 nm on Mar. 31, and he roughly estimates I ∼ 15. Clayton reports J = 14.5, K = 12.8 for Mar. 2. The star was not present in January, when a K observation of NGC 3504 was obtained with the CFHT Adaptive Optics Bonnette, and it is not present in other archival images, including JHK images taken in May 1996 at the Kitt Peak 2.1-m telescope (Elmegreen et al. 1997, A.J. 114, 1850) and in Dec. 1997 at Mt. Laguna Observatory. Note that this galaxy contains an apparent double nucleus, as well. P. Garnavich, Harvard-Smithsonian Center for Astrophysics, reports that inspection of a spectrum (range 840–900 nm) taken with the CFHT integral spectrograph OASIS on Mar. 31.312 (and provided by Emsellem) reveals what appears to be a broad Ca II P-Cyg profile characteristic of a supernova.

VARIABLE STAR NEAR NGC 4013? D. Bettoni, Osservatorio Astronomico di Padova, reports that further

examination of 12 Schmidt plates of NGC 4013 in the Asiago archive taken during 1983–1991 shows that the star noted on IAUC 6911 is present with no appreciable difference in brightness from the 1998 images. The star also appears at similar brightness on a May 1938 blue photograph present in the Carnegie Atlas of Galaxies (Sandage and Bedke 1994). B. F. Madore, NASA/IPAC Extragalactic Database, informed Bettoni that the NED has more than 100 galaxies (including NGC 4013) that are displayed with all stars subtracted from the field, without obvious mention of this fact. Furthermore, the Tully et al. images were converted to a logarithmic scale, whereby brighter objects may be depicted as fainter than on the original frames. Thus, though this star may be slightly variable, it does not appear to have an amplitude anywhere near 5 magnitudes.

1998 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green